### v5

# Randomized Approach to Matrix Completion: Applications in Collaborative Filtering and Image Inpainting

Abstract
We present a novel method for matrix completion, specifically designed for matrices where one dimension significantly exceeds the other. Our Columns Selected Matrix Completion (CSMC) method combines Column Subset Selection and Low-Rank Matrix Completion to efficiently reconstruct incomplete datasets. In each step, CSMC solves a convex optimization problem. We introduce two algorithms to implement CSMC, each tailored to problems of different sizes. A formal analysis is provided, outlining the necessary assumptions and the probability of obtaining a correct solution. To assess the impact of matrix size, rank, and the ratio of missing entries on solution quality and computation time, we conducted experiments on synthetic data. The method was also applied to two real-world problems: recommendation systems and image inpainting. Our results show that CSMC provides solutions of the same quality as state-of-the-art matrix completion algorithms based on convex optimization, while achieving significant reductions in computational runtime.
Keywords Matrix completion  $\cdot$ Column Subset Selection  $\cdot$ Low-rank models  $\cdot$ Image inpainting  $\cdot$ Collaborative filtering  $\cdot$ Nuclear Norm  $\cdot$ Convex Optimization
1 Introduction
Images, time series, and graphs are just a few examples of data types that can naturally be represented as real-valued matrices. In practice, these matrices often have missing entries due to data collection costs, sensor errors, or privacy considerations. To fully exploit the potential of matrix-based datasets, we need to address the missing data problem. Matrix completion (MC) methods leverage assumptions about underlying patterns in the data, which often manifest as a low-rank property of the matrix being analyzed. This low-rank characteristic reflects inherent correlations or regularities in the data, enabling effective recovery of missing entries.
Matrix completion has found applications in diverse fields such as signal processing, sensor localization, medicine and genomics [1, 2, 3, 4]. Its significance was underscored by the Netflix Prize challenge, which focused on predicting unknown ratings using collaborative filtering based on the preferences of many users [5, 6, 7]. The task involves completing the sparse user-movie rating matrix $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$, where $n_{1}$ (the number of users) is significantly larger than $n_{2}$ (the number of movies). This creates an imbalanced matrix in which most entries are missing, as users only rate a small subset of movies.
Singular Value Decomposition (SVD) forms the basis of many MC algorithms. It is commonly used to exploit the high correlation in the data by decomposing the matrix $\mathbf{M}$ as follows:
$\mathbf{M}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T},$ (1)
where $\mathbf{U}\in\mathbb{R}^{n_{1}\times n_{1}}$, $\mathbf{V}\in\mathbb{R}^{n_{2}\times n_{2}}$ are orthogonal, and $\mathbf{\Sigma}\in\mathbb{R}^{n_{1}\times n_{2}}$ is non-negative diagonal matrix of singular values, ordered according to their magnitude. The singular values indicate the strength of the variance along the corresponding directions, which are captured by the columns of $\mathbf{U}$ and $\mathbf{V}$. The rank of $\mathbf{M}$ is determined by the number of non-zero singular values. A common approach to MC is to identify the matrix of lowest rank that best approximates the observed entries. However, directly minimizing the rank is NP-hard [5, 8]. Instead, minimizing the nuclear norm, which is the sum of the singular values of the matrix, is often pursued [5]. Let $\Omega\subseteq\{1,\cdots,n_{1}\}\times\{1,\cdots n_{2}\}$ be the set of indices of the known entries, the convex optimization problem for the exact MC has the following form
[…]
where $\|\mathbf{X}\|_{*}$ denotes nuclear norm of the decision variable $\mathbf{X}$, $\mathcal{R}_{\Omega}:\mathbb{R}^{n_{1}\times n_{2}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}}$ sets all matrix elements to 00, except those in the $\Omega$ set, e.g.
$\mathcal{R}_{\Omega}(\mathbf{X})=\sum_{(i,j)\in\Omega}x_{ij}\mathbf{e}_{i}\mathbf{e}_{j}^{T},$ (3)
where $\mathbf{e}_{i}$ denotes standard basis vector with $1$ on the $i$-th coordinate and 00 elsewhere. Nuclear norm minimization is a widely adopted technique for low-rank matrix completion, known for its theoretical guarantees and empirical solid performance [9, 10, 5]. It can be solved in time polynomial time in the matrix dimension, using off-the-shelf Semidefinite Programming (SDP) solvers or tailored first-order methods [8, 10, 5]. However, in large-scale problems, solving these SDP programs, even with first-order methods, can be computationally challenging due to memory requirements [10].
This paper introduces a novel method for the low-rank matrix completion problem, which we call Columns Selected Matrix Completion (CSMC). The CSMC benefits from integrating two fields of modern linear algebra. In particular, it employs the Column Subset Selection (CSS) algorithm to reduce the size of the matrix completion problem. Recovering the most essential columns limits the computational costs of the matrix completion task. In the second stage, the recovered and previously known entries are used to recover $\mathbf{M}$ by solving the standard least squares problem [11, 12, 13]. CSMC is particularly useful when one dimension of $\mathbf{M}$ is much larger than the other, such as in the Netflix Prize scenario. The integration of the two approaches allows CSMC to decrease the computational cost of the convex matrix completion algorithms while maintaining their theoretical guarantees.
1.1 Notation and abbreviations
For a matrix $\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}}$ we denote by $\mathbf{x}_{i}$ its $i$-th column, and by $x_{ij}$ its $(i,j)$-th entry. Given the set of column indices $\mathcal{I}\in[n_{2}]$ we use $\mathbf{X}_{:,\mathcal{I}}$ to denote the column submatrix. We consider three types of matrix norms. The spectral norm of $\mathbf{X}$ is denoted by $\|\mathbf{X}\|_{2}$. The Frobenius and the nuclear norm are denoted by $\|\mathbf{X}\|_{F}$ and $\|\mathbf{X}\|_{*}$ respectively. Table 1 contains essential symbols and acronyms.
Notation Description
$\Omega$ set of indices of the observed entries
$\mathcal{R}_{\Omega}$ sampling operator (3)
$\mathcal{I}$ column indices for the column submatrix $\mathbf{C}$
$\mathbf{X}_{:,\mathcal{I}}$ column submatrix of $\mathbf{X}$ indicated by $\mathcal{I}$
$\mathbf{X}^{\dagger}$ Moore-Penrose inverse of $\mathbf{X}$
$\rho$ percentage of sampled entries
$\alpha$ percentage of sampled columns in the first stage
$r$ rank of $\mathbf{M}$
$\mu_{0}(\mathbf{M})$ the coherence parameter of $\mathbf{M}$
CSS Column Subset Selection
CSMC Columns Selected Matrix Completion
CSNN Columns Selected Nuclear Norm minimization
CSPGD Columns Selected Proximal Gradient Descent
NN Nuclear Norm minimization
PGD Proximal Gradient Descent
MF Matrix Factorization
Table 1: Notation and abbreviations
1.2 Outline of the paper
The rest of the paper is organized as follows: Section 2 covers related methods, our contribution, and the notations used. In Section 3, we present our method and provide formal analysis, particularly Theorem 3.2, which offers a lower bound on the probability of perfect reconstruction of the matrix, along with conditions on the minimum number of randomly selected columns and the minimum size of the sample of known elements of the reconstructed matrix that ensure achieving this estimate. Section 4 contains experimental results on synthetic and real datasets, confirming the effectiveness of the proposed method.
2 Related work and contribution
2.1 Low-rank matrix completion
The low-rank assumption is grounded in the observation that the real-world datasets inherently possess, or can be effectively approximated by, a low-rank matrix structure [10]. It plays a pivotal role in enhancing the efficiency of algorithms. The algorithms for the low-rank matrix completion typically leverage convex or non-convex optimization in Euclidean spaces [14, 9, 15, 16, 17, 18, 19], or employ Riemannian optimization [20, 21, 22]. The convex methods are the most mature and have strong and well-understood theoretical guarantees. The nuclear norm relaxation draws inspiration from the success of compressed sensing [10, 5, 23, 8] in which $\ell_{1}$ norm minimization is used to recover sparse signal instead of $\ell_{0}$ norm. The matrix rank may be expressed as the $\ell_{0}$ norm of the singular values vector $\mathbf{\sigma}$, while the nuclear norm is defined as its $\ell_{1}$ norm.
Another notable solution, alongside exact matrix completion (2), is presented by Mazumder et al. [15]. To address noisy data and prevent model overfitting, they consider a regularized optimization problem,
$\underset{\mathbf{X}}{\mathrm{minimize\ }}\frac{1}{2}\|\mathcal{R}_{\Omega}(\mathbf{X})-\mathcal{R}_{\Omega}(\mathbf{M})\|_{F}^{2}+\lambda\|\mathbf{X}\|_{*},$ (4)
where $\lambda$ is a regularization paramater. The problem (4) benefits from the closed form of the proximal operator for the objective function given by Singular Value Thresholding (SVT) and can be solved with Proximal Gradient descent (PGD) and its modifications [15].
Another class of convex optimization techniques focuses on minimizing the Frobenius norm error for the observed entries while applying constraints on the nuclear norm. It can be solved using the Frank-Wolfe (conditional-gradient descent) algorithm [24]. Despite its low computational cost per iteration, the Frank-Wolfe algorithm exhibits slow convergence rates. Consequently, numerous efforts have been undertaken to develop optimized variants of the algorithm [25, 26].
The non-convex optimization methods aim to minimize least squares error on the observed entries while constraining the rank of the solution with Matrix Factorization (MF) [27, 10],
$\min_{\mathbf{L}\in\mathbb{R}^{n_{1}\times k},\mathbf{R}\in\mathbb{R}^{n_{2}\times k}}\frac{1}{2}\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{L}\mathbf{R}^{T})\|^{2}_{F}.$ (5)
Chen and Chi [10] distinguish three classes of methods solving (5), Alternating Minimization, Gradient Descent and Singular Value Projection. For Alternating Minimization, Jain et al. [14] and Hardt [19] gave provable guarantees. However, these guarantees are worse in terms of coherence, rank and condition number of $\mathbf{M}$.
2.2 Column Subset Selection
Many commonly employed low-rank models, such as Principal Component Analysis (PCA), Singular Value Decomposition (SVD), and Nonnegative Matrix Factorization (NNMF) lack clear interpretability, leading to research on alternatives like CSS and CUR decomposition [28, 29, 30, 31, 32, 33]. CSS provides low-rank representation using the column submatrix of the data matrix $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$. Formally, given a matrix $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$ and a positive integer $d$, CSS seeks for the column submatrix $\mathbf{C}\in\mathbb{R}^{n_{1}\times d}$ such that
$\|\mathbf{M}-P_{C}(\mathbf{M})\|_{\xi}$ (6)
is minimized over all possible $\binom{n_{2}}{d}$ choices for the matrix $\mathbf{C}$. Parameter $\xi\in\{2,F\}$, where $\xi=2$ denotes the spectral and $\xi=F$ the Frobenius norm. Here,
$P_{C}=\mathbf{C}\mathbf{C}^{\dagger},$ (7)
is the projection onto the $d$-dimensional space spanned by the columns $\mathbf{C}$ and $\mathbf{C}^{\dagger}$ denotes the Moore-Penrose pseudoinverse [34]. On the other hand, CUR decomposition explicitly selects a subset of columns and rows from the original matrix to construct the factorization [31, 30].
Extensive research focuses on fully-observed matrices, which may not be practical for real-world applications [28, 29, 30, 31, 32, 33]. Bridging the research gap between CSS, CUR and matrix completion is a recently recognized challenge [35, 36, 37, 38]. In this context, two important questions emerge. The first one is about selecting meaningful columns (or rows) from partially observed data and applying CSS and CUR in matrix completion algorithms. Our method benefits from the fact that under standard assumptions about matrix incoherence, a high-quality column subset can be obtained by randomly sampling each column [39]. However, in the case of the coherent matrices, Wang and Singh [35] propose active sampling to dynamically adjust column selection during the sampling process, albeit with increased computational complexity.
The application of CUR and CSS in MC tasks typically involves altering the sampling scheme to relax assumptions about the matrix or reduce sample complexity [37, 38, 36]. However, estimating leverage scores, a key requirement in these methods, poses a significant challenge [35].
Cai et al. [3] showcase the potential of CUR in low-rank matrix completion, introducing a novel sampling strategy called Cross-Concentrating Sampling (CCS) and a fast non-convex Iterative CUR Completion (ICURC) algorithm. However, CCS requires $\mathcal{O}(r^{2}n\ln^{2}(n))$ observed entries, where $n=\max\{n_{1},n_{2}\}$, which is a factor of $r$ worse than the nuclear norm minimization-based approach. The optimization problem solved by ICURC imposes constraints on the rank of the matrix, thus introducing an additional parameter that requires tuning.
Our approach is related to the work of Xu et al., who presented CUR++, an algorithm for partially observed matrices [39]. It computes a low-rank approximation of the matrix based on randomly selected columns and rows, along with a subset of observed entries. However, the reconstruction procedure requires finding eigenvectors of the observed submatrices, which adds a computational burden.
2.3 Our contributions
This paper addresses the challenge of completing rectangular matrices by integrating principles from Column Subset Selection and Low-rank Matrix Completion. Our approach involves solving convex optimization problems. Our main contributions are as follows:
  1. 1.
    We introduce a two-staged matrix completion method, CSMC, dedicated to recovering rectangular matrices with one direction significantly larger than the other ($n_{1}<<n_{2}$).This structure is particularly well-suited for large-scale and asymmetrical data scenarios common in multi-source data integration tasks.
  2. 2.
    We offer a formal analysis, including Theorem 3.2, which establishes a lower bound on the probability of achieving perfect matrix reconstruction in the second stage of CSMC. This theorem contains necessary conditions regarding the minimum number of randomly selected columns ($d=\mathcal{O}(r\ln{r})$) and the minimum sample size of known elements ($|\Omega|=\mathcal{O}(n_{2}r\ln({n_{2}r}))$) required for accurate reconstruction of an incoherent matrix of rank $r$.
  3. 3.
    We develop two algorithms dedicated to problems of different sizes: Columns Selected Nuclear Norm (CSNN) and Columns Selected Proximal Gradient Descent (CSPGD).
  4. 4.
    We validate the CSMC method through numerical experiments on synthetic and real-world datasets, comparing its performance with matrix completion algorithms from the literature, specifically for recommendation systems and image inpainting.
3 Columns Selected Matrix Completion (CSMC) method overview
In the CSMC, the completion of the matrix $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$ is divided into two stages. In the first stage, CSMC selects the subset of $d$ columns and fills it with an established matrix completion algorithm. In the second stage, the least squares problem is solved. Each of the presented variants of the CSMC methods adopts the following procedure (Fig. 1).
Stage I: Sample and fill
CSMC selects $d$ columns of $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$ according to the uniform distribution over the set of all indices. Let $I\subseteq\{1,\cdots,n_{2}\}$ denote the set of the indices of the selected columns. The submatrix formed by selected columns, $\mathbf{C}:=\mathbf{M}_{:I}$, $\mathbf{C}\in\mathbb{R}^{n_{1}\times d}$ is recovered using the chosen matrix completion algorithm.
Stage II: Solve least squares
Let $\mathbf{\hat{C}}\in\mathbb{R}^{n_{1}\times d}$ be the output of the Stage I. The CSMC solves the following convex optimization problem,
$\min\limits_{\mathbf{Z}\in\mathbb{R}^{d\times n_{2}}}\frac{1}{2}\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\hat{C}}\mathbf{Z})\|_{F}^{2},$ (8)
where $\mathcal{R}_{\Omega}:\mathbb{R}^{n_{1}\times n_{2}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}}$ is the sampling operator (eq. 3). Let $\mathbf{\hat{Z}}\in\mathbb{R}^{d\times n_{2}}$ be a solution to the problem (8). The CSMC output with the matrix $\mathbf{\hat{M}}=\mathbf{\hat{C}}\mathbf{\hat{Z}}$.
Figure 1: The CSMC method overview.
Algorithm 1 CSMC-$\alpha$
Require: $\mathcal{R}_{\Omega}(\mathbf{M})$: $n_{1}\times n_{2}$-matrix
Require: $\alpha$: Ratio of the selected columns
$I\leftarrow\text{Uniformly sample }\alpha\cdot n_{2}$ indices
$\mathbf{C}_{\text{missing}}\leftarrow\mathcal{R}_{\Omega}(\mathbf{M})_{:I}$
$\mathbf{\hat{C}}\leftarrow\text{MC}(\mathbf{C}_{\text{missing}})$triangleright Complete submatrix $\mathbf{C}\in\mathbb{R}^{n_{1}\times d}$.
$\mathbf{\hat{Z}}\leftarrow\text{arg}\min\limits_{\mathbf{Z}\in\mathbb{R}^{n_{2}}}\frac{1}{2}\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\hat{C}}\mathbf{Z})\|_{F}^{2}$
$\mathbf{\hat{M}}\leftarrow\mathbf{\hat{C}}\mathbf{\hat{Z}}$
return $\mathbf{\hat{M}}$
3.1 Theoretical results
In their seminal work, Candes and Recht proved that, under certain assumptions, solving nuclear norm minimization leads to successful matrix recovery [17]. There has been a significant amount of research extending these results to noisy settings, improved bounds on the sample complexity and analyzing the robustness of the nuclear norm minimization algorithms to noise and outliers [40, 9, 18, 5]. Recht [9] greatly simplified the analysis and provided a foundation for understanding the potential and limitations of the method in practical matrix completion scenarios. The theoretical guarantees rely on properties such as matrix rank, number and distribution of observed entries and the fact that the singular vectors of the matrix are uncorrelated with the standard basis. The last property is formalized with the matrix coherence and defined as follows [17].
Definition 3.1.
Let $U$ be a subspace of $\mathbb{R}^{n}$ of dimension $r_{U}$ and $P_{U}$ be the orthogonal projection onto $U$. The coherence parameter of $U$ is defined as
$\mu(U)=\frac{n}{r_{U}}\max_{1\leq i\leq n}||P_{U}\mathbf{e}_{i}||_{2}^{2},$ (9)
where $\mathbf{e}_{i}$ for $i=1,\ldots n$ are the standard basis vectors of $\mathbb{R}^{n}$.
The coherence parameter of the rank-$\tilde{r}$ matrix $\mathbf{X}$ is given by $\mu_{0}(\mathbf{X})=\max\{\mu(U),\mu(V)\}$ where $U$ and $V$ are the linear spaces spanned by its $\tilde{r}$ left and right singular vectors respectively.
(a) $\mu(U)=1$.
(b) $21<\mu(U)<$ 2
(c) $\mu(U)=2$
Figure 2: Coherence of the one-dimensional subspace $U\subset\mathbb{R}^{2}$. 2(a) depicts $U$ spanned by $(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$ with the smallest coherence $\mu(U)=1$, while 2(c) presents $U$ spanned by one of the standard basis vectors $\mathbf{e}_{2}$ which achieves maximal coherence $\mu(U)=2$
We assume that observed entries are sampled uniformly at random. We begin with a discussion on how the crucial characteristics of the matrix $\mathbf{M}$ completion problem are transferred to the task of filling in the submatrix $\mathbf{C}$. Since $\mathbf{C}$ consists of randomly selected $d$ columns of $\mathbf{M}$, the rank of $\mathbf{C}$ is bounded by a minimum of $r$ and $d$, and the entries of $\mathbf{C}$ also follow a uniform distribution. The following theorem demonstrates that for an incoherent, well-conditioned matrix $\mathbf{M}$, if the submatrix $\mathbf{C}$ is composed of $\mathcal{O}(r\ln(rn_{2}))$ sampled columns, it will exhibit low coherence.
Theorem 3.1 (Corollary 3.6 in Cai et al. [41]).
Suppose that $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$ has rank $r$ and coherence bounded by $\mu_{0}(\mathbf{M})$. Suppose that $I\ \subseteq\{1,\ldots n_{2}\}$ is chosen by sampling uniformly without replacement to yield $\mathbf{C}=\mathbf{M}_{:I}$. Let $d$ be the number of sampled columns such that $d\geq 1.06\mu_{0}(\mathbf{M})r\ln(rn_{2})$. Then
$\mu_{0}(\mathbf{C})\leq 100\kappa^{2}({\mathbf{M}})\mu_{0}({\mathbf{M}}),$ (10)
with a probability at least $1-\frac{1}{n_{2}}$, where $\kappa(\mathbf{M})$ denotes the spectral condition number of $\mathbf{M}$, given by $\kappa(\mathbf{M}):=\|\mathbf{M}\|_{2}\|\mathbf{M}^{\dagger}\|_{2}$.
In this section, the primary focus lies on assessing the recovery capabilities of the least-squares problem during the second stage of the CSMC. In particular, we assume that submatrix $\mathbf{C}$ is fully observed or perfectly recovered and provide assumptions that solving
$\min_{\mathbf{Z}\in\mathbb{R}^{d\times n_{2}}}\frac{1}{2}\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{C}\mathbf{Z})\|_{F}^{2}$ (11)
will output $\mathbf{\hat{Z}}\in\mathbb{R}^{d\times n_{2}}$, such that $\mathbf{M}=\mathbf{C}\mathbf{\hat{Z}}$ holds with high probability. While similar to the CUR+ framework introduced in [39], our approach to solving (11) avoids the need to compute singular vectors of $\mathbf{C}$. The main result of this section is given by the Theorem 3.2. We believe that the following main result broadens the result of Xu et al. formulated as Theorem 2 in [39]. The proof of Theorem 3.2 is provided in the appendix.
Theorem 3.2.
Let $r$ be the rank of $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$ and let $\mathbf{C}\in\mathbb{R}^{n_{1}\times d}$ be a column submatrix of $\mathbf{M}$ formed by uniformly sampled without replacement $d$ columns. Let $\tilde{r}$ denote the rank of $\mathbf{C}$. Assume that for the parameter $\gamma>0$,
  1. 1.
    $d\geq 7\mu_{0}(\mathbf{M})r(\gamma+\ln r)$,
  2. 2.
    $|\Omega|\geq\tilde{r}n_{2}\mu_{0}(\mathbf{C})\left(\gamma+\ln{(\frac{n_{2}\tilde{r}}{2})}\right)$.
Let $\mathbf{\hat{Z}}\in\mathbb{R}^{d\times n_{2}}$ be the minimizer of the problem 11. Then, with a probability at least $1-3e^{-\gamma}$, we have $\mathbf{M}=\mathbf{C}\mathbf{\hat{Z}}$.
3.2 Implementation
The CSMC is a versatile method for solving matrix completion problems. It allows the incorporation of various matrix completion algorithms in the first step and least squares algorithms in the second step. This paper discusses two algorithms that implement the CSMC method, each tailored to address matrix completion problems of different sizes.
  1. 1.
    Columns Selected Nuclear Norm (CSNN) in which submatrix is filled with the exact nuclear norm minimization using SDP solver. The CSNN is dedicated to recovering the small and medium size $\mathbf{M}$. We found the first-order Splitting Conic Solver (SCS) as an efficient way to solve the SDP [42] in the Stage I. To find the optimum of the regression problem (8), we directly solve the least squares for the observed entries in each column. This approach is simple and allows to create fairly efficient distributed implementations.
  2. 2.
    Columns Selected Proximal Gradient Descent (CSPGD) in which inexact nuclear norm minimization (4) is solved by Proximal Gradient Descent (PGD). PGD is an efficient first-order method for convex optimization. It leverages the closed formula for the proximal operator for the objective function (4) and allows completing large matrices [15]. To find the optimum of the regression problem (8), we directly solve the least squares for the observed entries in each column.
We have developed and released open-source code implementing both the CSNN minimization algorithm and CSPGD methods. These algorithms support both Numpy arrays [43] and PyTorch tensors [44], with the latter offering the advantage of GPU acceleration for faster computation. The semidefinite programming is solved using the Splitting Conic Solver (SCS) [45]. The code is written in Python 3.10 and is available at https://github.com/ZAL-NASK/CSMC.
4 Numerical experiments
All testing examples were implemented in Python 3.10. For benchmarking purposes, the Matrix Factorization (MF) and Iterative SVD algorithms were utilized from the fancyimpute library [46].
4.1 Synthetic data set
To evaluate the proposed CSMC method, we compared the performance of the Columns Selected Nuclear Norm algorithm with the exact Nuclear Norm (NN) minimization in a controlled setting. The goal of each experiment was to recover a random $300\times 1000$ matrix $\mathbf{M}$ with the ratio of missing entries $\rho$. To control the rank of $\mathbf{M}$ equal to $r$, the test matrix was generated as a product of the $n_{1}\times r$ matrix $\mathbf{A}$ and $r\times n_{2}$ matrix $\mathbf{B}$. Matrices $\mathbf{A}$ and $\mathbf{B}$ were generated in two steps. In the first one, matrix entries were sampled from the normal distribution $\mathcal{N}(0,1)$. In the second step, noise matrices with the ratio $0.3$ of non-zero entries were added to each matrix.
Experiment S I
To assess the sample complexity of the presented methods, we varied the ratio of known entries $\rho$ and the number of sampled columns $d$. Every setup was executed over $N_{\text{trial}}=20$ independent trial runs. The performance of the algorithms NN and CSNN-${\alpha}$, $\alpha=\{0.1,\cdots,0.9\}$ was evaluated over all trials. To inquire how many columns should be sampled to achieve a satisfying algorithm relative error,
$\epsilon:=\frac{\|\mathbf{M}-\mathbf{\mathbf{\hat{M}}}\|_{F}}{\|\mathbf{M}\|_{F}},$ (12)
we calculated an empirical cumulative distribution function (ECDF) to display the proportion of trials achieved given approximation error [47, 48]. ECDF is defined as $\hat{F}_{S}:\mathbb{R}_{+}\rightarrow[0,1]$,
$\hat{F}_{S}(a)=\frac{|\{s|\epsilon_{s}\leq a\}|}{|S|},$ (13)
where each trial is represented as $s$, $\epsilon_{s}$ denotes the relative error reached by the solution of $s$, and $S$ denotes the set of all trials for the given parameters setup. We also compared the analyzed algorithms’ runtime (in seconds). Since CSMC methods consist of two stages, while other algorithms are one staged, we did not compare iteration numbers of algorithms to achieve solutions of prescribed quality.
Experiment S II
With the same experiment settings we benchmarked CSNN-$\alpha$ for $\alpha\in\{0.1,\cdots,0.5\}$ algorithm, Matrix Factorization (MF) [49, 27], Iterative SVD [50], and Nuclear Norm minimization (NN).
Results
Experiments were executed on the Linux workstation equipped with 11th Gen Intel(R) Core(TM) i7-1165G7 @ 2.80GHz (8 cores) and 32GB RAM.
Fig. 3 shows that sampling CSNN-$0.2$ offered a tenfold (10x) increase in time savings compared to the NN algorithm and maintains a solution quality. Although MF was faster than CSNN-$\alpha$ algorithms (Fig. 5), the magnitude of the relative error for CSNN-$0.3$ and CSNN-$0.5$ was significantly smaller for rank-5 and rank-10 cases, respectively. In that case, both algorithms returned a solution of superior quality (Fig. 4).
Figure 3: Results S I: ECDF (13) and runtimes depending on the matrix rank and rate of the known entries for NN and CSNN-$\alpha$, $\alpha\in\{0.1,\cdots,0.9\}$, $\mathbf{M}\in\mathbb{R}^{300\times 1000}$. Sampling with CSNN-$0.2$ resulted in a tenfold (10x) improvement in time efficiency over the NN algorithm while preserving solution quality.
Figure 4: Results S II: ECDF (13) curves depending on the matrix rank and rate of the known entries for NN and CSNN-$\alpha$, $\alpha\in\{0.1,\cdots,0.9\}$, $\mathbf{M}\in\mathbb{R}^{300\times 1000}$. The relative error magnitude for CSNN-$0.3$ and CSNN-$0.5$ was notably lower in rank-5 and rank-10 scenarios, respectively, compared to MF.
Figure 5: Results S II: ECDF (13) curves depending on the matrix rank and rate of the known entries for NN and CSNN-$\alpha$, $\alpha\in\{0.1,\cdots,0.9\}$, $\mathbf{M}\in\mathbb{R}^{300\times 1000}$.
4.2 Recommendation system
Here, we assessed the performance of CSNN and CSPGD algorithms within the context of the Netflix Prize scenario, where matrix completion was employed as a rating prediction technique. Data sets used for the benchmarking are publicly available data sets from the Movielens research project [51].
Settings
To benchmark CSNN and CSPGD algorithms, we constructed two data sets: Movie Lens Small and Movie Lens Big. The Movie Lens Small was represented by $\mathbf{M}\in\mathbb{R}^{140\times 668}$ matrix obtained from the Movie Lens Small dataset, which contained 100 000 5-star ratings applied to 9742 movies by 610 users. Since the original matrix was too large for SDP solvers, we followed the procedure described in [3] to obtain a submatrix of the desired size. Specifically, we sorted the data by user frequency rate and took data containing rates made by the top 60% users. Then, we sorted the obtained data by movie frequency rate and selected the top 50% movies. The obtained matrix had $\rho=0.25$ known entries. We evaluate CSPGD algorithms on the Movielens 25M data set, containing 25 million ratings applied to 62,000 movies by 162,000 users. Again, due to the large size and low observation rate, we extracted a $654\times 27813$ matrix $\mathbf{M}$. The known entries rate $\rho$ was equal to 0.09.
We assessed the performance of the NN and CSNN-$\alpha$ for $\alpha\in\{0.3,0.4,0.5,0.7\}$ and CSPGD-$\alpha$ for $\alpha\in\{0.3,0.5\}$. We followed previous work [3] and employed the Cross-Validation method. In each trial of the experiment, $\Omega$ set was randomly split into training and testing sets denoted by $\Omega_{\text{train}}$ and $\Omega_{\text{test}}$ [52]. Specifically, we randomly selected $\rho$ rate of the observed set, and by assigning the null values to the rest of the entries, we constructed $\mathcal{R}_{\Omega_{\text{train}}}(\mathbf{M})$ matrix. We evaluated each algorithm on the $\Omega_{\text{test}}$ set. We conducted 20 independent experimental trials under each scenario.
We compared Normalized Mean Absolute Error (NMAE) calculated as
$\text{NMAE}=\frac{1}{|\Omega_{\text{test}}|(m_{\text{max}}-m_{\text{min}})}\sum_{(i,j)\in\Omega_{\text{test}}}|\mathbf{\hat{M}}-\mathbf{M}|,$ (14)
where $m_{\text{max}}$ and $m_{\text{min}}$ denote the maximum and minimum rating, respectively, and $\Omega_{\text{test}}$ denotes the set of indices in test set. This metric is widely used to assess collaborative filtering tasks [3]. The quality of the recommendation was also measured as hit-rate, defined as
$\text{HR}=\frac{\#hits}{|\Omega_{\text{test}}|},$ (15)
where a predicted rating was considered a hit if its rounded value equals the actual rating in the test set [3].
Results
Fig. 6 shows that the predictions of CSNN-$0.7$ maintained the quality of the NN algorithm in terms of NMAE: (0.16 vs 0.13) and HR (0.23 vs 0.25), and results in runtime savings from 57 seconds to 40 seconds (Table 2). CSPGD algorithms were much faster than PGD (Fig. 7). The HR values for PGD and CSPGD-$0.3$ were equal to 0.28 and 0.24, respectively (Table 3).
NMAE HR Time [s]
mean std mean std mean std
Algorithm
CSNN-0.3 0.271 0.028 0.172 0.005 10.137 1.118
CSNN-0.4 0.240 0.060 0.191 0.005 17.677 2.005
CSNN-0.5 0.248 0.176 0.208 0.006 28.673 2.419
CSNN-0.7 0.156 0.011 0.232 0.004 39.893 8.595
NN 0.127 0.001 0.255 0.006 56.924 2.435
Table 2: Results on Movie Lens Small dataset; NN and CSNN-$\alpha$ algorithms, $\alpha\in\{0.3,0.4,0.5,0.7\}$.
Figure 6: Results on Movie Lens Small dataset; NN and CSNN-$\alpha$ algorithms, $\alpha\in\{0.3,0.4,0.5,0.7\}$.
NMAE HR Time [s]
mean std mean std mean std
Algorithm
CSPGD-0.3 0.138 0.001 0.245 0.002 23.032 0.321
CSPGD-0.5 0.135 0.001 0.252 0.001 27.774 0.299
PGD 0.119 0.001 0.281 0.001 449.026 9.502
Table 3: Results on Movie Lens Big dataset; PGD and CSPGD-$\alpha$ algorithms for $\alpha\in\{0.3,0.5\}$.
Figure 7: Results on Movie Lens Big dataset; PGD and CSPGD-$\alpha$ algorithms for $\alpha\in\{0.3,0.5\}$.
4.3 Image recovery
Image inpainting, a technique in image processing and computer vision, fills in missing or corrupted parts. Additionally, estimating random missing pixels can be treated as a denoising method or utilized to accelerate rendering processes. Low-rank models are highly effective for this task, leveraging the assumption that images typically exhibit low-rank structures. The primary information of the image matrix is dominated by its largest singular values while setting the smallest singular values to zero can be done without losing essential details.
Settings
The data set contained ten grey-scaled pictures of bridges downloaded from the public repository 111image source (https://pxhere.com/pl/). Images were represented as $240\times 360$ matrices. We assessed the performance of the algorithms among 100 independent trials (ten trials per picture). The quality of the reconstructed image was assessed with the signal-to-noise ratio (SNR)
$\text{SNR}(\mathbf{\hat{M}})=20\log_{10}\left(\frac{\|\mathbf{M}\|_{F}}{\|\mathbf{\hat{M}}-\mathbf{M}\|_{F}}\right)$ (16)
and the relative error (12).
Results
The CSNN-$0.7$ solutions maintained the quality of the NN in terms of SNR (Table 4). As shown in Fig. 9, CSNN-$0.5$ offered considerable time savings and good relative error. Fig. 8 presents one of the completed pictures.
Bridge restored with NN.
Bridge restored with CSNN-0.7
Figure 8: Image inpainting; NN and CSNN-0.7 algorithms, $\rho=20\%$ known entries.
Figure 9: Results for image inpainting; NN and CSNN-$\alpha$ algorithms for $\alpha\in\{0.3,0.4,0.5,0.7\}$.
SNR Relative error ($\epsilon$) Time [s]
Known entries ratio $\rho$ 0.4 0.2 0.1 0.4 0.2 0.1 0.4 0.2 0.1
Algorithm
CSNN_0.3 18.844 10.829 7.366 0.114 0.289 0.438 16.306 12.026 19.365
CSNN_0.4 17.708 12.245 11.491 0.130 0.245 0.268 24.151 17.997 28.123
CSNN_0.5 17.851 15.245 14.172 0.128 0.173 0.196 35.719 24.424 38.647
CSNN_0.7 21.134 18.526 16.458 0.088 0.119 0.151 73.670 44.948 65.639
NN 25.221 21.377 18.582 0.055 0.085 0.118 113.215 83.196 88.966
Table 4: Results for image inpainting; NN and CSNN-$\alpha$, $\alpha\in\{0.3,0.4,0.5,0.7\}$.
5 Conclusion
This paper introduces the Columns Selected Matrix Completion (CSMC) method, which enhances the time efficiency of nuclear norm minimization algorithms for low-rank matrix completion. CSMC employs a two-stage approach: first, it applies a low-rank matrix completion algorithm to a reduced problem size, followed by minimizing the least squares error. Theoretical analysis provides provable guarantees for perfect matrix recovery under standard assumptions.
Numerical simulations show that CSMC significantly reduces computational time while maintaining high accuracy in the reconstructed data. Compared to alternative methods, CSMC demonstrates competitive performance in approximating missing entries. This enhanced time efficiency is particularly valuable for tasks such as recommendation systems, image processing, and other matrix-based applications, where large, incomplete datasets must be effectively recovered. Reducing computational time while minimizing accuracy loss is crucial for enabling efficient processing of incomplete data, which is central to many real-world applications in pattern recognition.
CSMC relies on uniform column sampling and nuclear norm-based algorithms, but its flexible framework allows for the integration of alternative sampling strategies and matrix completion techniques. Future work will explore adaptive sampling approaches, such as those proposed in recent studies [37, 35], to handle high-coherence matrices and reduce sample complexity.
Acknowledgments
The authors thank Dr. Aikaterini Aretaki for her critical review and comments on the manuscript.
Appendix A Proof of Theorem 3.2 and supporting theorems
We present a comprehensive proof of Theorem 3.2, accompanied by the development of two additional theorems, namely A.2 and A.4, which are our original contributions. The proofs for these supplementary theorems are provided within this work. Additionally, the cited theorems from the literature have their proofs outlined therein.
We assume that $\mathbf{C}\in\mathbb{R}^{n_{1}\times d}$ is a column submatrix of $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$. $\mathbf{C}$ is obtained by sampling uniformly without replacement $d$ columns of $\mathbf{M}$. Let $r$ be the rank of $\mathbf{M}$ and let $\tilde{r}$ be the rank of $\mathbf{C}$. We assume that the compact SVD of $\mathbf{C}$ is given by,
$\mathbf{C}=\mathbf{\tilde{U}}\mathbf{\tilde{\Sigma}}\mathbf{\tilde{V}^{T}},$ (17)
where $\mathbf{\tilde{U}}\in\mathbb{R}^{n_{1}\times\tilde{r}}$, $\mathbf{\tilde{V}}\in\mathbb{R}^{n_{2}\times\tilde{r}}$, and $\mathbf{\tilde{\Sigma}}\in\mathbb{R}^{\tilde{r}\times\tilde{r}}$.
We denote as $\mu_{0}(\mathbf{M})$ and $\mu_{0}(\mathbf{C})$ the coherence parameters of $\mathbf{M}$ and $\mathbf{C}$.
We denote as $\tilde{U}$ the subspace spanned by the first $\tilde{r}$ left singular vectors of $\mathbf{C}$. The orthogonal projection onto $\tilde{U}$ is given by
$P_{\tilde{U}}=\mathbf{\tilde{U}}\mathbf{\tilde{U}}^{T}.$ (18)
The following lemma can be found in [39] and shows that under the incoherence assumption, uniform sampling outputs a high-quality solution to the CSS problem, with high quality.
Theorem A.1 (Theorem 9 in Xu et al. [39]).
Let $\tilde{U}_{\psi}$ denote subspace spanned by the $\psi$ first left singular vectors of $\mathbf{C}$ and let $P_{\tilde{U}_{\psi}}$ denote orthogonal projection on $\tilde{U}_{\psi}$. Then for parameter $\gamma>0$, with probability $1-2e^{-\gamma}$ we have
$\|\mathbf{M}-P_{\tilde{U}_{\psi}}\mathbf{M}\|_{2}^{2}\leq\sigma_{\psi+1}^{2}\left(1+\frac{2n_{2}}{d}\right),$ (19)
if $d\geq 7\mu(U_{\psi})\psi(\gamma+\ln{\psi})$, where $d$ denotes number of sampled columns in $\mathbf{C}$, $\sigma_{\psi+1}$ is the $(\psi+1)$-th largest singular value of $\mathbf{M}$ and $\mu(U_{\psi})$ is the coherence of the subspace spanned by $\psi$ first left singular vectors of $\mathbf{M}$,
The following remark is the immediate consequence of the Theorem A.1.
Remark A.1.
Let $\psi\geq r$, where $r$ is the rank of $\mathbf{M}$. Then $\sigma_{\psi+1}=\sigma_{r+1}=0$ and with probability $1-2e^{-\gamma}$,
$\|\mathbf{M}-P_{\tilde{U}}\mathbf{M}\|^{2}_{2}=0,$ (20)
provided that $d\geq 7\mu_{0}(\mathbf{M})r(\gamma+\ln{r})$.
We now bound the distance measured in the spectral norm between $\mathbf{M}$ and $\mathbf{\hat{M}}=\mathbf{C}\mathbf{\hat{Z}}$. To do that, we assume that the objective function $g:\mathbb{R}^{\tilde{r}\times n_{2}}\rightarrow\mathbb{R}$,
$g(\mathbf{Y})=\frac{1}{2}\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{Y})\|_{F}^{2}$ (21)
is strongly convex [53], where the strong convexity is defined as follows.
Definition A.1 (Boyd and Vandenberghe [53]).
A function $g:\mathbb{R}^{\tilde{r}\times n_{2}}\rightarrow\mathbb{R}$ is strongly convex with parameter $\beta>0$ if $h(\mathbf{X})=g(\mathbf{X})-\frac{\beta}{2}\|\mathbf{X}\|_{F}^{2}$ is convex.
The following remarks provide a helpful characterization of strongly convex functions.
Remark A.2 (Boyd and Vandenberghe [53]).
A function $g:\mathbb{R}^{\tilde{r}\times n_{2}}\rightarrow\mathbb{R}$ is strongly convex with parameter $\beta>0$ if $g$ is everywhere differentiable and
$g(\mathbf{Y})\geq g(\mathbf{X})+\langle\nabla g(\mathbf{X}),\mathbf{Y}-\mathbf{X}\rangle+\frac{\beta}{2}\|\mathbf{Y}-\mathbf{X}\|_{F}^{2},$ (22)
for any $\mathbf{X},\mathbf{Y}\in\mathbb{R}^{\tilde{r}\times n_{2}}$, where inner product is defined as
$\langle\mathbf{X},\mathbf{Y}\rangle=\text{tr}(\mathbf{X}^{T}\mathbf{Y}).$ (23)
Remark A.3 (Boyd and Vandenberghe [53]).
If $g$ is twice differentiable, then $g(\mathbf{X})$ is strongly convex with parameter $\beta>0$ if $\nabla^{2}g(\mathbf{X})\succeq\beta\mathbf{I}$ for any $\mathbf{X}\in\mathbb{R}^{\tilde{r}\times n_{2}}$.
We now provide the formula for the first-order and the second-order partial derivatives of $g$. Let $y_{sw}$ denote the $(s,w)$ entry of the matrix $\mathbf{Y}$, and $m_{lw}$ denote $(l,w)$ entry of $\mathbf{M}$ then the partial derivative of $g$ with respect to $y_{sw}$ is given by following formula,
$\frac{\partial g}{\partial y_{sw}}=-\sum\limits_{l:(l,w)\in\Omega}(m_{lw}-\sum_{i=1}^{\tilde{r}}{\tilde{u}_{li}}y_{iw})\tilde{u}_{ls}=\sum\limits_{l:(l,w)\in\Omega}(\sum_{i=1}^{\tilde{r}}{\tilde{u}_{li}}y_{iw}-m_{lw})\tilde{u}_{ls}.$ (24)
The second-order partial derivative with respect to $y_{pq}$ and $y_{sw}$ is equal to
$\frac{\partial^{2}g}{\partial y_{sw}\partial y_{pq}}=\begin{cases}\sum\limits_{l:(l,w)\in\Omega}\tilde{u}_{ls}\tilde{u}_{lp}&w=q,\\ 0&w\neq q.\end{cases}$ (25)
The following theorem shows that under certain assumptions, $\mathbf{\hat{M}}=\mathbf{C}\mathbf{\hat{Z}}$ is close to the matrix $\mathbf{M}$ in the spectral norm. Those assumptions include that i) $P_{\tilde{U}}\mathbf{M}$ is close to $\mathbf{M}$, ii) $g$ is strongly convex with parameter $\beta$. The following proofs take inspiration from the work of Xu et al. [39].
Theorem A.2.
Suppose that $\|\mathbf{M}-P_{\tilde{U}}\mathbf{M}\|_{2}^{2}\leq\Delta$ for the parameter $\Delta>0$, and let $\tilde{U}$ be the subspace of $\mathbb{R}^{n_{1}}$ spanned by the first $\tilde{r}$ left singular vectors of $\mathbf{C}$, $P_{\tilde{U}}=\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}$. Assume that $g$ is strongly convex with a parameter $\beta$, then
$\|\mathbf{M}-\mathbf{\hat{M}}\|_{2}^{2}\leq\Delta+\frac{2n_{1}\tilde{r}\Delta}{\beta}.$ (26)
Proof.
Set $\mathbf{Y}_{1}=\mathbf{\mathbf{\tilde{U}}^{T}}\mathbf{M}$ , then from the assumptions
$\|\mathbf{M}-\mathbf{\tilde{U}}\mathbf{Y}_{1}\|_{2}^{2}\leq\Delta.$ (27)
Thus
$\frac{1}{r^{\prime}}\|\mathbf{M}-\mathbf{\tilde{U}}\mathbf{Y}_{1}\|_{F}^{2}\leq\|\mathbf{M}-\mathbf{\tilde{U}}\mathbf{Y}_{1}\|_{2}^{2}\leq\Delta,$ (28)
where $r^{\prime}$ denotes the rank of the matrix $(\mathbf{M}-\mathbf{\tilde{U}}\mathbf{Y}_{1})\in\mathbb{R}^{n_{1}\times n_{2}}$ and
$\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{Y}_{1})\|_{F}^{2}\leq\|\mathbf{M}-\mathbf{\tilde{U}}\mathbf{Y}_{1}\|_{2}^{2}\leq r^{\prime}\Delta.$ (29)
Since any of the matrix dimensions bounds rank, $r^{\prime}\leq n_{1}$ (for thick matrices $n_{1}<n_{2}$)
$\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{Y}_{1})\|_{F}^{2}\leq n_{1}\Delta.$ (30)
Let $\mathbf{\hat{Z}}$ be a solution to the problem 11, then
$f(\mathbf{\hat{Z}})=\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{C}\mathbf{\hat{Z}})\|_{F}^{2}=\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{\tilde{\Sigma}}\mathbf{\tilde{V}}^{T}\mathbf{\hat{Z}})\|_{F}^{2}$ (31)
and $\mathbf{\hat{Y}}:=\mathbf{\tilde{\Sigma}}\mathbf{\tilde{V}}^{T}\mathbf{\hat{Z}}$ must be a minimum of $g$. Indeed, assume that $g(\mathbf{Y}_{2})<g(\mathbf{\hat{Y}})$ for some $\mathbf{Y}_{2}\neq\mathbf{\hat{Y}}$, i.e.
$\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{C}\mathbf{\hat{Z}})\|_{F}^{2}$ $=\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{\hat{Y}})\|_{F}^{2}$ (32)
$>\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{Y_{2}})\|_{F}^{2}$
$=\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{C}\underbrace{\mathbf{\tilde{V}}\mathbf{\tilde{\Sigma}^{-1}}\mathbf{Y_{2}}}_{\mathbf{X}_{2}})\|_{F}^{2}$
and $f(\mathbf{X}_{2})<f(\mathbf{\hat{Z}})$ where $f$ is the objective function in problem 11.
We bound distance between $\mathbf{Y}_{1}$ and $\mathbf{\hat{Y}}$ using strong convexity of $g$. Since $\mathbf{\hat{Y}}$ is minimum of $g(\mathbf{Y})$ (21), then $\nabla g(\mathbf{\hat{Y}})=\mathbf{0}$, and
$\langle\nabla g(\mathbf{\hat{Y}}),\mathbf{Y}_{1}-\mathbf{\hat{Y}}\rangle=0.$ (33)
Thus, by the Remark A.2,
$g(\mathbf{Y}_{1})\geq g(\mathbf{\hat{Y}})+\frac{\beta}{2}\|\mathbf{Y}_{1}-\mathbf{\hat{Y}}\|_{F}^{2}.$ (34)
Using the definition of $g$ given in (21),
$\frac{\beta}{2}\|\mathbf{Y}_{1}-\mathbf{\hat{Y}})\|_{F}^{2}$ $\leq\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{Y}_{1})\|_{F}^{2}-\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{\hat{Y}})|_{F}^{2}$ (35)
$\leq\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{Y}_{1})\|_{F}^{2}.$
Combining (35) and (30),
$\|\mathbf{Y}_{1}-\mathbf{\hat{Y}}\|_{F}^{2}\leq\frac{2}{\beta}\|\mathcal{R}_{\Omega}(\mathbf{M})-\mathcal{R}_{\Omega}(\mathbf{\tilde{U}}\mathbf{Y}_{1})\|_{F}^{2}\leq\frac{2n_{1}}{\beta}\Delta.$ (36)
Using the triangle inequality and the fact that Frobenius norm of a matrix is always greater than or equal to its spectral norm,
$\|\mathbf{M}-\mathbf{\hat{M}}\|_{2}^{2}$ $=\|\mathbf{M}-\mathbf{C}\mathbf{\hat{Z}}\|_{2}^{2}$ (37)
$\leq\|\mathbf{M}-\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}\|_{2}^{2}+\|\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}-\mathbf{C}\mathbf{\hat{Z}}\|_{2}^{2}$
$\leq\|\mathbf{M}-P_{\tilde{U}}\mathbf{M}\|_{2}^{2}+\|\mathbf{\tilde{U}}\mathbf{\tilde{U}^{T}}\mathbf{M}-\mathbf{C}\mathbf{\hat{Z}}\|_{F}^{2}.$
The first component on the right side is bounded by the assumptions,
$\|\mathbf{M}-P_{\tilde{U}}\mathbf{M}\|_{2}^{2}\leq\Delta.$ (38)
To bound the second component, we use the definition of the $\mathbf{Y}_{1}$ and $\mathbf{\hat{Y}}$, and the fact that the product of the Frobenius norms bounds Frobenius norm of the product of the two matrices.
$\|\mathbf{\tilde{U}}\mathbf{\tilde{U}}^{T}\mathbf{M}-\mathbf{\tilde{U}}\mathbf{\tilde{\Sigma}}\mathbf{\tilde{V}}^{T}\mathbf{\hat{Z}}\|_{F}^{2}$ $=\|\mathbf{\tilde{U}}\mathbf{Y}_{1}-\mathbf{\tilde{U}}\mathbf{\hat{Y}}\|_{F}^{2}$ (39)
$\leq\|\mathbf{\tilde{U}}\|_{F}^{2}\|\mathbf{Y}_{1}-\mathbf{\hat{Y}}\|_{F}^{2}$
Using (36) and the fact that $\|\mathbf{\tilde{U}}\|_{F}=\sqrt{\hat{r}}$, since columns of $\mathbf{\tilde{U}}\in\mathbb{R}^{n_{1}\times\tilde{r}}$ are orthonormal.
$\|\mathbf{\tilde{U}}\|_{F}^{2}\|\mathbf{Y}_{1}-\mathbf{\hat{Y}}\|_{F}^{2}\leq\frac{2n_{1}\tilde{r}\Delta}{\beta},$ (40)
implying
$\|\mathbf{M}-\mathbf{\hat{M}}\|_{F}^{2}\leq\Delta+\frac{2n_{1}\tilde{r}\Delta}{\beta}.$ (41)
To bound the parameter $\beta$ of the strong convexity, we use Remark A.3 and bound the smallest eigenvalue of the Hessian of $g$. Following [39], to do that, we use the following result of Tropp [54].
Theorem A.3 (Theorem 5 in Xu et al. [39] derived from Theorem 2.2 in Tropp [54] ).
Let $\mathcal{X}$ be a finite set of the positive-semidefinite (PSD) matrices with dimension $k\times k$, and suppose that
$\max_{\mathbf{X}\in\mathcal{X}}\lambda_{\max}(\mathbf{X})\leq B$ (42)
for some parameter $B>0$, where $\lambda_{\max}(\mathbf{X})$ is the maximum eigenvalue of $\mathbf{X}$. Sample $\{\mathbf{X}_{1},\ldots,\mathbf{X}_{\Psi}\}$ uniformly at randommo from $\mathcal{X}$ without replacement. Compute:
$\mu_{\min}:=\Psi\lambda_{\min}(\mathbb{E}\mathbf{X}_{1}),$ (43)
and
$\mu_{\max}:=\Psi\lambda_{\max}(\mathbb{E}\mathbf{X}_{1}),$ (44)
where $\mathbb{E}\mathbf{X}_{1}$ is the expected value of a random variable $\mathbf{X}_{1}$, $\lambda_{\max}(\mathbb{E}\mathbf{X}_{1})$ and $\lambda_{\min}(\mathbb{E}\mathbf{X}_{1})$ denote its maximum and minimum eigenvalue.
$P\big{(}\lambda_{\max}(\sum_{j=1}^{\Psi}\mathbf{X}_{j})\geq(1+\rho)\mu_{\max}\big{)}\leq k\exp{\frac{-\mu_{\max}}{B}}[(1+\rho)\ln{(1+\rho)}-\rho],$ (45)
for parameter $\rho\in[0,1)$.
$P\big{(}\lambda_{\min}(\sum_{j=1}^{\Psi}\mathbf{X}_{j})\leq(1-\rho)\mu_{\min}\big{)}\leq k\exp{\frac{-\mu_{\min}}{B}}[(1-\rho)\ln{(1-\rho)}+\rho],$ (46)
for parameter $\rho\geq 0$.
Theorem A.4.
Let $\gamma>0$ be a parameter. With probability $1-e^{-\gamma}$ we have that the objective function $g:\mathbb{R}^{\tilde{r}\times n_{2}}\rightarrow\mathbb{R}$ defined in (21) is strongly convex with parameter $\beta>0$, provided that
$|\Omega|\geq\tilde{r}n_{2}\mu(\tilde{U})\left(\gamma+\ln{(\frac{n_{2}\tilde{r}}{2})}\right).$ (47)
Proof.
By remark A.3 to bound strong convexity, we can instead bound the smallest eigenvalue of the Hessian matrix $\mathbf{H}=\nabla^{2}g$.
The Hessian matrix of a function $g$ is a $\tilde{r}n_{2}\times\tilde{r}n_{2}$ matrix. Let us assume that second-order derivative with respect to the $y_{sw}$ and $y_{pq}$ entries of matrix $\mathbf{Y}$ is the $(\tilde{r}(s-1)+w$, $\tilde{r}(p-1)+q)$ entry of the Hessian matrix. Then using eq. 25 the Hessian matrix $\mathbf{H}$ can be written as
$\begin{pmatrix}\mathbf{H}^{1,1}&\mathbf{H}^{1,2}&\ldots&\mathbf{H}^{1,{\tilde{r}}}\\ \mathbf{H}^{2,1}&\ddots&\ldots&\mathbf{H}^{2,{\tilde{r}}}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{H}^{{\tilde{r}},1}&\mathbf{H}^{{\tilde{r}},2}&\ldots&\mathbf{H}^{{\tilde{r}},\tilde{r}}\\ \end{pmatrix}$ (48)
where $\mathbf{H}^{s,q}$ is a diagonal $n_{2}\times n_{2}$ matrix containing partial derivatives $\frac{\partial^{2}g}{\partial y_{sw}\partial y_{pq}}$ for $w,q\in\{1,\ldots,n_{2}\}$.
$\mathbf{H}^{s,q}=\begin{pmatrix}\sum\limits_{l:(l,1)\in\Omega}\tilde{u}_{ls}\tilde{u}_{lp}&0&\ldots&0\\ 0&\ddots&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\sum\limits_{l:(l,n_{2})\in\Omega}\tilde{u}_{ls}\tilde{u}_{lp}\\ \end{pmatrix}$ (49)
Then the Hessian of $g$ is a sum of the random matrices,
$\mathbf{H}=\sum\limits_{(i,j)\in\Omega}\mathbf{\tilde{u}}_{i:}\mathbf{\tilde{u}}_{i:}^{T}\otimes\mathbf{e}_{j}\mathbf{e}_{j}^{T},$ (50)
where $\mathbf{e}_{j}\in\mathbb{R}^{n_{2}}$ is a standard basis vector and $\mathbf{\tilde{u}}_{i:}\in\mathbb{R}^{r}$ is a vector defined by the $i$-th row of matrix $\mathbf{\tilde{U}}$, and $\otimes$ denote the Kronecker product.
Thus the Hessian of $g$ is a sum of the $|\Omega|$ random matrices of the form
$\mathbf{H}^{i,j}:=\mathbf{\tilde{u}}_{i:}\mathbf{\tilde{u}}_{i:}^{T}\otimes\mathbf{e}_{j}\mathbf{e}_{j}^{T},$ (51)
where $\mathbf{\tilde{u}}_{i:}\mathbf{\tilde{u}}_{i:}^{T}\in\mathbb{R}^{\tilde{r}\times\tilde{r}}$ is PSD and $\mathbf{e}_{j}\mathbf{e}_{j}^{T}\in\mathbb{R}^{n_{2}\times n_{2}}$ is also PSD. Thus,
$\mathbf{H}=\sum\limits_{(i,j)\in\Omega}\mathbf{H}^{i,j}.$ (52)
Each $\mathbf{H}^{i,j}$ is PSD as the Kronecker product of the two PSD matrices. Moreover,
$\lambda_{\max}(\mathbf{H}^{i,j})$ $=\lambda_{\max}(\mathbf{\tilde{u}}_{i:}\mathbf{\tilde{u}}_{i:}^{T})$ (53)
$\leq\max_{1\leq j\leq r}|u_{ij}|^{2}$
$\leq\frac{\tilde{r}\mu(\tilde{U})}{n_{1}},$
for each $i=1,\cdots n_{1}$, $j=1,\cdots n_{2}$.
Let $(i_{1},j_{1})$ be the first pair of indices in $\Omega$, i.e., the indices of the first observed entry in $\mathbf{M}$. The expected value of the random matrix $\mathbf{H}^{i_{1},j_{1}}$ is given by
$\mathbb{E}(\mathbf{H}^{i_{1},j_{1}})$ $=\frac{1}{n_{1}n_{2}}\sum\limits_{l=1}^{n_{1}}\sum\limits_{q=1}^{n_{2}}\mathbf{H}^{l,q}$ (54)
$=\frac{1}{n_{1}n_{2}}\mathbf{\tilde{U}}^{T}\mathbf{\tilde{U}}\otimes\mathbf{I}_{n_{1}\times n_{1}}$
$=\frac{1}{n_{1}n_{2}}\mathbf{I}_{\tilde{r}\times\tilde{r}}\otimes\mathbf{I}_{n_{1}\times n_{1}}$
$=\frac{1}{n_{1}n_{2}}\mathbf{I}_{\tilde{r}n_{1}\times\tilde{r}n_{1}}.$
Following the notation of Theorem A.3,
$\mu_{\min}=|\Omega|\lambda_{\min}(\mathbb{E}(\mathbf{H}^{i_{1},j_{1}}))=\frac{|\Omega|}{n_{1}n_{2}},$ (55)
and
$\max_{ij}\lambda_{\max}(\mathbf{H}^{i,j})\leq\frac{\tilde{r}\mu(\tilde{U})}{n_{1}}=B.$ (56)
Combining Theorem A.3 with $\rho=\frac{1}{2}$ and eq. (52)
$P(\lambda_{\min}(\mathbf{H})\leq\frac{1}{2}\mu_{\min})\leq\tilde{r}n_{2}\exp{\left(\frac{-\mu_{\min}}{B}\right)}\left(\frac{1-\ln{2}}{2}\right)\leq\frac{\tilde{r}n_{2}}{2}\exp{\left(\frac{-\mu_{\min}}{B}\right)},$ (57)
implying,
$P\left(\lambda_{\min}(\mathbf{H})\leq\frac{|\Omega|}{2n_{1}n_{2}}\right)\leq\frac{\tilde{r}n_{2}}{2}\exp{\left(-\frac{|\Omega|}{\tilde{r}n_{2}\mu(\tilde{U})}\right)},$ (58)
Hence, with probability at least $1-e^{-\gamma}$,
$\lambda_{\min}(\mathbf{H})\geq\frac{|\Omega|}{2n_{1}n_{2}},$ (59)
provided that
$|\Omega|\geq\tilde{r}n_{2}\mu(\tilde{U})\left(\gamma+\ln{(\frac{n_{2}\tilde{r}}{2})}\right).$ (60)
Theorem 3.2 can be proved by combining the results of Theorems A.2 and A.4.
Proof of Theorem 3.2.
Since $d\geq 7\mu_{0}(\mathbf{M})r(\gamma+\ln r)$, from the Remark A.1, we have
$\|\mathbf{M}-P_{\tilde{U}}\mathbf{M}\|_{2}^{2}=0,$ (61)
with the probability at least $1-2e^{-\gamma}$.
From Theorem A.4, the fact that $\mu_{0}(\mathbf{C})>\mu(\tilde{U})$ and the fact $|\Omega|\geq\tilde{r}n_{2}\mu_{0}(\mathbf{C})\left(\gamma+\ln{(\frac{n_{2}\tilde{r}}{2})}\right)$, function $g$ is $\beta$-strongly convex with probability at least $1-e^{-\gamma}$.
The probability the fact, that eq. (61) holds and $g$ is $\beta$-strongly convex is grater or equal than $(1-3e^{-\gamma})+(1-e^{-\gamma})-1$.
Thus we can apply Theorem A.2 with $\Delta=0$ and show that
$\|\mathbf{M}-\mathbf{\hat{M}}\|_{2}^{2}\leq 0,$ (62)
implying $\mathbf{M}=\mathbf{\hat{M}}$ with probability at least $1-3e^{-\gamma}$.